OpenStudy (anonymous):

A bacteria culture begins with 4 bacteria which double in size every hour. How many bacteria exist in the culture after 8 hours?

OpenStudy (anonymous):

64 bacteria

OpenStudy (anonymous):

@Isaiah.Feynman what do you think

OpenStudy (anonymous):

no wait is 32

OpenStudy (isaiah.feynman):

There is a basic formula for exponenetial growth I found on wikipedia.

OpenStudy (raden):

1st h = 4 * 2 = 8 2nd h = 8 * 2 = 16 .... so on, just times 2

OpenStudy (isaiah.feynman):

\[x(t)=a \times b^{\frac{ \tau }{ t }}\]

OpenStudy (isaiah.feynman):

Use that formula.

OpenStudy (anonymous):

how do i use it

OpenStudy (isaiah.feynman):

Okay wait, I noticed I made a mistake the exponent is t/tau, NOT tau/t.

OpenStudy (anonymous):

ok so what do i plug in them

OpenStudy (isaiah.feynman):

t=8hours (convert to minutes!), tau is 1 hour(convert to minutes!), b=2( the bacteria DOUBLES), and a=4

OpenStudy (anonymous):

1024? right?

OpenStudy (isaiah.feynman):

I didn't calculate it, but should be right. Who would make a mistake with such easy formula